OpenStudy (anonymous):

3.A yogurt shop offers 6 different flavors of frozen yogurt and 12 different toppings. How many choices are possible for a single serving of frozen yogurt with one topping? 144 72 36 665,280 4.In how many different orders can you line up 8 cards on a table? 8 1 1,680 40,320 5.There are 10 students participating in a spelling bee. In how many ways can the students who go first and second in the bee be chosen? 1 way 90 ways 3,628,800 ways 45 ways

OpenStudy (anonymous):

all by the counting principle \(6\times 12\) for the first one

OpenStudy (anonymous):

\[8!=8\times 7\times 6\times 5\times 4\times 3\times 2\] for the second

OpenStudy (anonymous):

and finally \(10\times 9\) for the third

OpenStudy (anonymous):

@satellite73 For the third one, why did you only times 10 by 9? I thought there was more to that. I used the combinations formula for this problem and got 45 ways.

OpenStudy (anonymous):